OpenStudy (anonymous):

Is the function y = x2 - 6 even, odd, or neither? A. even B. odd C. neither

OpenStudy (anonymous):

even

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

y = x^2 - 6 is the same as y = x^2 - 6x^0 notice how all the exponents are even numbers so because you have this polynomial with all even number exponents, the function is even